OpenStudy (anonymous):

Inverse of f(x) = 3x - 5

OpenStudy (anonymous):

Make f(x)= y...so you have y=3x-5...now swap the x and y so you have x=3y-5..solve for y...that is how to do inverse functions

OpenStudy (anonymous):

I got all of that, but how do I solve for y? That's the part I'm stuck on @nelsonjedi

OpenStudy (anonymous):

x=3y-5...first add 5 to both sides of the equation..what do you have now?

OpenStudy (anonymous):

5x = 3y

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

adding 5 to x does not give you 5x does it?

OpenStudy (anonymous):

sorry 5 + x

OpenStudy (anonymous):

or x + 5 @nelsonjedi

OpenStudy (anonymous):

Correct now you have x+5=3y...now how to you convert 3y to y?

OpenStudy (anonymous):

divide by 3 @nelsonjedi

OpenStudy (anonymous):

Corrrect and what ever you do to one side of the equation you do to the other. So divide both sides by 3 and what do you have?

OpenStudy (anonymous):

x + 5/3 = y?

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

\[\frac{ x+5 }{ 3 }=y\]

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

Welcome